A bullet is found embedded in the wall of a room 3.2 m above the floor. The bullet entered the wall going upward at an angle of 34.3°. How far from the wall was the bullet fired if the gun was held 1.5 m above the floor?


1
Expert's answer
2022-02-22T13:50:14-0500

"tan34.3^0=\frac{3.2-1.5}{d}."

"d=\frac{1.7}{tan34.3^0}=2.49m."